OpenStudy (acabral058):

Solve the system using any method (provide steps please) \sqrt(3x)-\sqrt(2y)=2\sqrt(3) 2\sqrt(2x)+\sqrt(3y)=\sqrt(2)

OpenStudy (kevin):

I don't understand the question, can u upload the picture?

OpenStudy (acabral058):

@kevin no i can't screenshot from my computer. I have to solve the system using substitution or elimination. Its square root of 3x minus square root of 2y is equal to 2 square root 3. The second equation is 2 square root of 2x plus square root of 3y is equal to sqaure root too.

OpenStudy (acabral058):

*square root 2 not too sorry

alones (alones):

Like this? \[\sqrt(3x)-\sqrt(2y)=2\sqrt(3)\] \[\ 2\sqrt(2x)+\sqrt(3y)=\sqrt(2)\]

OpenStudy (acabral058):

@ AloneS yes

alones (alones):

Alright lol, so yo need help with them..or something? adn @alones like this LOL

OpenStudy (acabral058):

@AloneS yes. I have to use substitution or elimination.

alones (alones):

Ahha aright.

alones (alones):

waitt so those two problems come together, or they're separate?

OpenStudy (acabral058):

together. I am solving a system.

zepdrix (zepdrix):

\[\large\rm \color{royalblue}{\sqrt{3x}-\sqrt{2y}=2\sqrt3}\]Let's try to line up the x's. We'll multiply this first equation by 2sqrt2,\[\large\rm \color{royalblue}{2\sqrt{6x}-4\sqrt{y}=4\sqrt6}\]Then,\[\large\rm \color{orangered}{2\sqrt{2x}+\sqrt{3y}=\sqrt2}\]Multiplying our second equation by sqrt3 gives us,\[\large\rm \color{orangered}{2\sqrt{6x}+3\sqrt{y}=\sqrt6}\]Now the x's match, yay! So you can subtract one from the other, and solve for y.